# Unifying Demonstration Selection and Compression for In-Context Learning

### v1

@@ 4 Experiment
In this section, we comprehensively evaluate UniICL over out-domain datasets to verify its effectiveness. Additionally, we detailedly introduce of involved datasets and evaluation metrics in Appendiex C.
@@ 4.1 Baselines
Naive Vicuna-7b serves as the fundamental baseline fed into actual demonstrations. AutoCompressor recurrently compresses demonstrations into virtual tokens. We employ their Llama2-7b version [^6] https://github.com/princeton-nlp/AutoCompressors.. LLMLingua is a coarse-to-fine demonstration pruning method based on dropping uninformative words. We employ their released 7b version [^7] https://github.com/microsoft/LLMLingua., of which compressor is a fine-tuned Llama-2. For a meaningful comparison, we replace target LLMs of LLMLingua (GPT-3.5-Turbo or Claude-v1.3) with the Vicuna-7b. ICAE [^8] https://github.com/getao/icae. compresses demonstrations into soft prompts via a LoRA-adapted Llama2-7b, which is most similar to UniICL in model architecture and compression mode. Additionally, since selection augmentation is involved in the training of UniICL, we utilize the popular Sentence-BERT (S-BERT) Reimers and Gurevych (2019) as the dense retriever to construct an ICL pipeline for the above methods, serving as simple but effective selection-based baselines.
@@ 4.2 Settings
Considering the involved datasets and computation efficiency, we set the max allowed input length limit to 1,024 for both compression and generation. For a fair comparison, we set the allowed window of baselines to 1,024 and their compression ratio to 12. We fix the learning rate to 8e-5 and use Adam as the optimizer, and the effective batch size is 32 (8 GPUs data parallelism and 4 steps gradient accumulation). Additionally, we conducted all experiments on 8*NVIDIA A5000 24G GPUs based on BFloat 16 data type, and we set the evaluated shot to 8 for understanding tasks and 5 for generative tasks for illustration. Except for the MS MARCO preprocessed by Liang Wang et al. (2022)  [^9] https://github.com/microsoft/unilm/tree/master/simlm., the other evaluation datasets are publicly achieved at hugging face.
We apply S-BERT to pre-rank and output the top 10 similar candidates from training sets according to each inference input and employ UniICL to perform selection among them in practice due to computation efficiency for high-resource ICL. On the contrary, the low-resource ICL setting fixes the candidate demonstrations to 20 for all inference inputs, performing pre-ranking and selecting as well.
@@ 4.3 Results
Since the virtual tokens naturally summarize semantic information of preceding sequences, we evaluate UniICL on the out-domain MS MARCO dataset in Table 3. UniICL significantly outperforms the sparse retrieval method BM25 algorithm and other compression methods. Notably, we don’t compare UniICL with the popular retrieval models Reimers and Gurevych (2019); Wang et al. (2024) since most of them are fine-tuned on this dataset, unfair for comparison.
@@ 5.1 Compression Ratio
During training, the compression ratio is dynamically sampled from 2 to 16. We mix up 2,000 instances from the in-domain validation set, 1,000 for XSUM, and 1,000 for CICERO to select the compression ratio for UniICL in Figure 7, with the backbone of Vicuna and BlueLM respectively. Specifically, UniICL compresses the latter cut-off part while keeping the former ones uncompressed. Therefore, we can measure the dense information quality of the same content with different compression ratios by ROUGE-1 since it is more sensitive to token-level differences. The performance is relative smoothing when the compression ratio changes from 4×4$\times$4 × to 12×12$\times$12 ×. However, when it comes to 16×16$\times$16 ×, an obvious drop occurs. Therefore, we set the compression ratio to 12 by default and apply this ratio to all experiments. The 1024×1024$\times$1024 × compression ratio is equal to compressing anything to a single virtual token, due to the maximum allowed input length for compression being 1,024.
@@ 5.2 Efficiency Analysis
In UniICL, we incorporate an additional 17M trainable parameters into the 7b backbone, accounting for an approximate increase of 2.4%. We additionally evaluate the memory costs inference latency of UniICL and other compression methods in Figure 1. Notably, compressing each demonstration independently and allowing horizon concatenation among different virtual demonstrations convert the extra compression latency be a one-shot deal that the virtual tokens can be stored, formulating a Demonstration Bank (DB) for further reusing. In this case, UniICL will eliminate the extra latency if the selected demonstrations have been compressed and cached (UniICL+Caching). Despite this, parallel computation facilitates the compressing process, resulting in minimal throughput degradation (UniICL and Baseline). What’s more, the naive 7B LLM occurs memory explosion for 8-shot settings, while UniICL successfully scales up to 64-shots within 24GB CUDA allocation.
5.3 Necessarity of UniICL
Intuitively, UniICL applies a “heavy” LLM as the compressor, but the generator and the compressor share all parameters without loading double parameters as LLMLingua, bypassing the OoD problem between the compressor and generator. Additionally, ICAE switches the LoRA parameters for compression and generations respectively, which appear to make the compressor and the generator share parameters. However, the compressor and the generator in ICAE are still two different LLMs inherently after assembling LoRA weight. Furthermore, the continual switch also breaks the latency-free merit of LoRA, resulting in more inference latency than the projection layer in UniICL, as depicted in Figure 1.
@@ Appendix A In-Domain Evaluation
Method XSUM CICERO R-1 R-2 R-L R-1 R-2 R-L Vicuna 19.9 5.0 13.5 17.3 3.3 14.3    +UniICL 27.2 9.2 22.5 24.6 5.1 21.7
Table 4: The in-domain results of XSUM and CICERO. Vicuna generates responses conditioned on the truncated inference input and UniICL first compresses the over-length input into virtual tokens.
We conduct the zero-shot in-domain generation evaluation on the entire test set of XSUM and CICERO in Table 4, by compressing the latter half to virtual tokens and keeping the former unmodified. UniICL significantly outperforms the baseline Vicuna-7b, indicating the compressed virtual tokens can provide the original truncated information by recovering the cut-off parts after supervised fine-tuning.
@@ Appendix B Results on BlueLM
Method #-shots XSUM ARXIV R-1 R-2 R-L R-1 R-2 R-L BlueLM 0-shot 23.5 6.8 17.6 30.9 7.7 24.7 1-shot 19.1 4.8 12.1 23.0 3.6 19.0 UniICL 1-shot 24.0 6.9 18.0 31.4 7.7 25.2 2-shot 25.0 7.3 18.8 30.8 7.3 24.8 5-shot 25.0 7.3 18.8 30.8 7.3 24.8 UniICL 1-shot 25.2 7.4 18.9 31.6 7.9 25.4 2-shot 25.4 7.6 19.1 31.9 8.0 25.6 5-shot 26.5 7.9 20.3 32.1 8.0 25.5
@@ C.1 Datasets
Dataset # words (96,512] (512,1024] (1024,1536] XSUM - 10,000 4,697 CICERO 10,000 - - XSUM (Ctr) 5,000

### v2

@@ 4.1 Baselines
Naive Vicuna-7b serves as the fundamental baseline fed into actual demonstrations. AutoCompressor recurrently compresses demonstrations into virtual tokens. We employ their Llama2-7b version [^5] https://github.com/princeton-nlp/AutoCompressors.. LLMLingua is a coarse-to-fine demonstration pruning method based on dropping uninformative words. We employ their released 7b version [^6] https://github.com/microsoft/LLMLingua., of which compressor is a fine-tuned Llama-2. For a meaningful comparison, we replace target LLMs of LLMLingua (GPT-3.5-Turbo or Claude-v1.3) with the Vicuna-7b. ICAE [^7] https://github.com/getao/icae. compresses demonstrations into soft prompts via a LoRA-adapted Llama2-7b.
Additionally, since selection augmentation is involved in the training of UniICL, we utilize the popular Sentence-BERT (S-BERT) Reimers and Gurevych (2019) as the dense retriever to construct an ICL pipeline for the above methods, serving as simple but effective selection-based baselines.
@@ 4.2 Settings
Considering the involved datasets and computation efficiency, we set the max allowed input length limit to 512 for both compression and generation. For a fair comparison, we set the allowed window of baselines to 512 and their compression ratio to the same as UniICL. We fix the learning rate to 8e-5 and use Adam as the optimizer, and the effective batch size is 32 (8 GPUs data parallelism and 4 steps gradient accumulation). Additionally, we conducted all experiments on 8*NVIDIA A5000 24G GPUs based on BFloat 16 data type, and we set the evaluated shot to 8 for understanding tasks and 5 for generative tasks for illustration because of marginal ICL gains and memory costs.
We apply S-BERT to pre-rank and output the top 10 similar candidates from training sets according to each inference input for all baselines. UniICL is employed to perform selection among them in practice due to computation efficiency for high-resource ICL. On the contrary, the low-resource ICL setting fixes the random candidate demonstrations to 20 for all inference inputs, performing pre-ranking and selecting as well.
@@ 4.3 Results
Since the virtual tokens naturally summarize semantic information of preceding sequences, we evaluate UniICL on the out-domain MS MARCO dataset in Table 3. UniICL significantly outperforms the sparse retrieval method BM25 algorithm and other compression methods. Notably, we don’t compare UniICL with the popular retrieval models Reimers and Gurevych (2019); Wang et al. (2024) since most of them are fine-tuned on this dataset, which is unfair for comparison.
@@ 5.1 Compression Ratio
During training, the compression ratio is dynamically sampled from 2 to 16. We mix up 2,000 instances from the in-domain validation set, 1,000 for XSUM, and 1,000 for CICERO to select the compression ratio for UniICL in Figure 7, with the backbone of Vicuna and BlueLM respectively. Specifically, UniICL compresses the latter cut-off part while keeping the former ones uncompressed. Therefore, we can measure the dense information quality of the same content with different compression ratios by ROUGE-1 since it is more sensitive to token-level differences. The performance is relative smoothing when the compression ratio changes from 4×4$\times$4 × to 12×12$\times$12 ×. However, when it comes to 16×16$\times$16 ×, an obvious drop occurs. Therefore, we set the compression ratio to 12 by default and apply this ratio to all experiments. The 512×512$\times$512 × compression ratio is equal to compressing anything to a single virtual token, due to the maximum allowed input length for compression being 512.
Figure 8: The efficiency comparison between UniICL and other compression methods in CoLA with the number of shots increasing from 0 to 64. Memory explodes are represented as *, corresponding to the break of the line chart.
Method GPUHours TFLOPs TMACs Vicuna 1.5 86,20 4,309 Vicuna-1k 1.9 31,664 15,832 UniICL 1.6 22,437 11,218
Table 4: The computation efficiency of UniICL.
@@ 5.2 Efficiency Analysis
In UniICL, we incorporate an additional 17M trainable parameters into the 7b backbone, accounting for an approximate increase of 0.24%. We additionally evaluate the memory costs inference latency of UniICL and other compression methods in Figure 8. With the help of the Demonstration Bank (DB), UniICL will eliminate the extra latency if the selected demonstrations have been compressed and cached (UniICL+Caching). Despite this, parallel computation facilitates the compressing process, resulting in minimal throughput degradation (UniICL and Baseline). What’s more, the naive 7B LLM occurs memory explosion for 8-shot settings and other compression methods perform up to 32-shot, while UniICL successfully scales up to 64-shots within 24GB CUDA allocation.
Additionally, We demonstrate the inference computation and GPU hours in Table 4, by using 1,024 random legal tokens as inputs and forcing models to generate 128 tokens. Notably, UniICL (without DB) compresses the former half, and the latter half is fed into the generator directly, while Vicuna and Vicuna-1k are distinguished in window limitations. Results indicate that minimal GPU hours increased due to the parallel computation of forward, although the extra compression of UniICL surges the computation. Additionally, Vicuna with a 1k window limitation surges both GPU hours and TFLOPs because long input brings significant computation and latency in generation.
5.3 Training Deviation
To quantify the performance gains brought by the learnable projection layer. We tune Vicuna and BlueLM with comparable parameters (17M) with LoRA in Table 5, setting the rank to 32. UniICL still outperforms LoRA-adapted LLMs with a 512 window limitation, indicating that the truncation indeed brings performance degradation.
@@ Appendix A In-Domain Evaluation
Backbone Method XSUM CICERO R-1 R-2 R-L R-1 R-2 R-L Vicuna-7b Vicuna 19.9 5.0 13.5 17.3 3.3 14.3    +LoRA 25.4 7.5 17.3 28.1 10.5 25.6 Vicuna-1k 27.3 8.7 19.7 30.5 11.3 27.4    +LoRA 31.2 11.0 23.1 34.1 13.5 30.2 UniICL 30.0 10.2 22.3 32.6 12.2 28.8 BlueLM-7b BlueLM 15.0 3.6 10.4 17.6 3.1 15.0    +LoRA 23.1 7.6 17.4 21.9 7.8 19.8 BlueLM-1k 28.1 9.9 22.8 25.1 9.2 23.1    +LoRA 30.8 10.5 24.6 31.2 10.8 27.4 UniICL 30.4 10.2 23.7 29.2 10.0 26.6
Table 5: The in-domain results and ablation studies on XSUM and CICERO. 1k represents the extending 1k window limitation, while others have a limitation of 512.
We conduct the zero-shot in-domain generation evaluation on the entire test set of XSUM and CICERO in Table 5 by compressing the latter half to virtual tokens and keeping the former unmodified. UniICL significantly outperforms the baselines, indicating the compressed virtual tokens can provide the original truncated information by recovering the cut-off parts after supervised fine-tuning. Although extending the window to 1k, Vicuna and BlueLM still underperform UniICL, indicating that compressed virtual tokens filter noise information to some extent.
@@ Appendix B Results on BlueLM
Method #-shots XSUM ARXIV R-1 R-2 R-L R-1 R-2 R-L BlueLM 0-shot 15.0 3.6 10.4 30.9 7.7 24.7 1-shot 19.1 4.8 12.1 23.0 3.6 19.0 UniICL 1-shot 24.0 6.9 18.0 31.4 7.7 25.2 2-shot 25.0 7.3 18.8 30.8 7.3 24.8 5-shot 25.0 7.3 18.8 31.9 7.8 26.0 UniICL 1-shot 25.2 7.4 18.9 31.6 7.9 25.4 2-shot 25.4 7.6 19.1 31.9 8.0 25.6 5-shot 26.5 7.9 20.3 32.1 8.0 25.5
@@ C.1 Datasets
Dataset # words (96,512] (512,1024] (1024,1536] XSUM - 10,000 4,697 CICERO 10,000 - - SUPER-NI - 10,000 7,000 XSUM (Ctr) 5,000